# From Label Error Detection to Correction: A Modular Framework and Benchmark for Object Detection Datasets

### v1

@@ Abstract
Object detection has advanced rapidly in recent years, driven by increasingly large and diverse datasets. However, label errors, defined as missing labels, incorrect classification or inaccurate localization, often compromise the quality of these datasets. This can have a significant impact on the outcomes of training and benchmark evaluations. Although several methods now exist for detecting label errors in object detection datasets, they are typically validated only on synthetic benchmarks or limited manual inspection. How to correct such errors systemically and at scale therefore remains an open problem. We introduce a semi-automated framework for label-error correction called RECD (Rechecked). Building on existing detectors, the framework pairs their error proposals with lightweight, crowd-sourced microtasks. These tasks enable multiple annotators to independently verify each candidate bounding box, and their responses are aggregated to estimate ambiguity and improve label quality. To demonstrate the effectiveness of RECD, we apply it to the class pedestrian in the KITTI dataset. Our crowdsourced review yields high-quality corrected annotations, which indicate a rate of at least 24% of missing and inaccurate annotations in original annotations. This validated set will be released as a new real-world benchmark for label error detection and correction. We show that current label error detection methods, when combined with our correction framework, can recover hundreds of errors in the time it would take a human to annotate bounding boxes from scratch. However, even the best methods still miss up to 66% of the true errors and with low quality labels introduce more errors than they find. This highlights the urgent need for further research, now enabled by our released benchmark.
@@ I Introduction
Deep neural networks (DNNs) have become state-of-the-art for extracting information from large-scale visual data [1, 2, 3, 4, 5, 6]. Training them for specialized and safety-critical domains, such as autonomous driving [1, 2, 5] or medical imaging [3, 6, 7, 8], requires vast amounts of high-quality annotated data [9, 10, 11, 12]. Crowdsourcing offers a cost- and time-effective solution for building large-scale vision datasets [13] and enables scalable annotation [14, 15, 16], particularly when tasks are broken down into simple, quick interactions known as microtasks [14, 13]. The quality of labels is critical because label errors such as missing labels, incorrect classes, or poor localization degrade model performance, generalization capability [17, 18, 19, 20, 21, 22, 23, 24] and benchmark validity [25, 17]. Although recent research has proposed automated methods for detecting such errors in image classification [25, 26, 27], semantic segmentation [28], and object detection [29], few works have focused on correcting them [26, 30, 31, 32] and no use error detection method to reduce the reannotation costs.
We propose RECD (see Figure 1), a semi-automated framework that transitions from label error detection to correction for object detection datasets. RECD combines automated label error proposals from existing methods [33, 34, 35] with lightweight validation via microtasks. Aggregated responses capture both label correctness and ambiguity, enabling scalable and cost-effective correction. We apply RECD to the pedestrian class in the KITTI dataset [36]. Pedestrians in KITTI represent a high-ambiguity, high-error-rate class that is both practically important (e.g. in autonomous driving) and methodologically challenging due to size variation, occlusion, and scene clutter [37]. Our annotations reveal at least 24.6% of previously unlabeled or incorrectly located pedestrian labels. These corrected annotations will be made publicly available as a new benchmark for evaluating methods for detecting and correcting label errors. Our analysis shows that current detection methods can be rechecked faster than humans can draw the bounding boxes for labeling. However, up to 66% of errors remain undetected, highlighting the need for improved detection strategies, which are now supported by our benchmark.
@@
We propose RECD a scalable, cost-efficient framework that utilizes state-of-the-art label error detection methods, extending them to label error correction in object detection datasets which can find hundreds of errors faster than humans can annotate the bounding boxes.
We introduce ambiguity-aware labeling through soft labels yielding high quality object labels. We demonstrate that lower quality data might introduce more errors than it is fixing. We release corrected pedestrian annotations for KITTI, enabling evaluation of label error detection methods on real-world label noise.
With our framework we detected at least 24.6% of missing and inaccurate pedestrian annotations. However, we also find up to 66% of the errors can currently not be found by error correction methods, highlighting the importance of this research now enabled by our released benchmark.
We make our source code publicly available at https://github.com/JonathanKlees/rechecked .
@@ II Related Work
II-A Identifying Label Errors in Object Detection Datasets.
@@
In recent years, there has been a research focus on label error detection in object detection. Cleanlab’s tool ObjectLab [35] assesses the quality of labels in object detection datasets by comparing predicted bounding boxes with ground truth, and provides a score describing the quality of the labels. Another method, based on the instance-wise object detection loss [34], uses a two-stage object detector to detect real label errors by monitoring the regression and classification losses and using the discrepancy between predictions and ground truth labels. MetaDetect [33] is a method that performs meta classification and meta regression for predictions. Potential label errors are those predicted boxes that are evaluated as false positive ( with ground truth), but the meta classifier estimates a high probability of being correct. The method LidarMetaDetect [38, 39] is a further development of MetaDetect, as it detects label errors in 3D object detection datasets. Bär et al. [40] developed a label refinement network that improves noisy localization labels. While we build on this recent research we go one step further and aim at correcting these labels instead of just detecting them.
II-B Label Error Correction.
@@
In addition to detecting label errors, correcting them has also become an important step in improving label quality. Bernhardt et al. [41] developed an active label cleaning framework to prioritize samples for reannotation by comparing estimated label correctness and label difficulty. CROWDLAB [42] is another tool of Cleanlab that classifies crowdsourced labels for refinement to identify which samples require re-annotation. In the work of Ma et al. [31], annotators was formed to re-annotate the COCO [43] and Google Open Images [44] datasets to compare model performance on original and re-annotated data. Kim et al. [32] combines foundation models with label corrections provided by human annotators to publish PASCAL+, an updated version of the Pascal VOC dataset for semantic segmentation and Yun et al. [45] applied their re-labeling strategy, ReLabel, to re-annotate the ImageNet training set and published a corrected version. Our work focuses on the process of label error correction based on error detection, with the aim of reducing annotation costs.
II-C Ambiguity and Soft Labels in Computer Vision Datasets.
@@
In common computer vision benchmark datasets, images or objects are usually associated with a single hard class label. However, recent research shows that this concept is not representative of the ambiguity in real data [46, 47, 48, 49, 50, 51, 52, 53, 54, 23] and even harms model performance [20, 21, 22, 55, 23, 24]. In contrast to this, soft or probabilistic labels [56] indicate a probability distribution across several classes reflecting that the correct label for an instance is not always straightforward and multiple annotators may disagree on the label. Soft labels in object detection datasets are rare but there exist methods that simulate soft labels based on single-class annotations through deep learning [57, 58, 59]. We extend this research by showing that the label quality is essential for correcting labels and poor quality data might even introduce more errors than it corrects. Furthermore, we release these soft labels as benchmark for future error detection and correction research.
@@ III Method RECD
The most challenging task in label error detection is detecting missing labels, i.e., objects present in the image that are missing in the ground truth (GT) annotations. Unlike inaccurate bounding box localizations and incorrect class labels, which only require verifying existing annotations and are therefore easier to identify, detecting missing labels requires a complete review of each image. We propose a semi-automatic framework that combines label error detection methods with a microtask based review process to efficiently correct label errors. It consists of three main stages: object detection, label error proposal, and microtask based correction. We provide an overview of the pipeline in Figure 2 and explain the interaction between automated and human components to detect missing labels and bounding box annotations with an inaccurate localization.
III-A Object Detection Models
In the initial phase of the proposed framework, a trained object detector is used to obtain predictions on the respective dataset establishing the foundation of our framework. We use YOLOX [60], a one-stage detector, and Cascade R-CNN [61], a two-stage detector. YOLOX is an anchor-free detector with a decoupled head and applies stronger data augmentations, such as Mosaic or MixUp, to enhance the network’s performance. The Cascade R-CNN comprises of a sequence of detectors in its multi-stage architecture. These detectors are trained sequentially, with the output of one detector serving as the training set for the subsequent one. This particular architecture is motivated by the objective of mitigating both overfitting and quality mismatch during inference.
III-B Label Error Detection Methods
The second stage of our framework automatically generates proposals for label errors, which are then to be reviewed manually in the third stage. These proposals are generated by integrating existing label error detection methods, including MetaDetect [33], loss-based instance-wise scoring [34] and ObjectLab [35]. Each method predicts a score that reflects the likelihood of a predicted box indicating a label error for every detected bounding box. The considered methods rely on the predictions of an object detector, which are compared to the ground truth annotations to identify overlooked objects or misfitting bounding boxes.
III-C Error Correction
From the previous stage, we obtain a collection of label error proposals that need to be manually evaluated. We created short and easy-to-answer questions consisting of a single visual task, e.g. whether the highlighted object is a human or not as illustrated in Figure 3. These kinds of micro-questions are also referred to as microtasks [62]. With microtasks, annotators can quickly answer one simple question at a time. We repeat the task with several people at the same time resulting in an estimate of the underlying distribution of the outcomes associated to this question.
More complex questions, such as ’What is a pedestrian?’, are broken down into multiple microtasks, such as ’Is it a human?’ and ’What is the person doing?’. See subsection IV-B for the microtasks used to construct the validated annotations of the KITTI dataset. We can use the calculated soft label probability to determine whether a proposed label error is indeed erroneous or not.
@@ IV-A Dataset Construction
We applied our label error detection and correction framework to the KITTI [36] 2D object detection dataset. We focus on the pedestrian class in KITTI as a challenging and safety-relevant use case. We consider a single class and dataset to enable a thorough investigation of the label error correction process, ensuring high-quality validation and detailed analysis, while the overall design of our framework remains dataset- and class-agnostic to support future generalization. The dataset contains 7,481 training images and 7,518 test images showing street scenes. Since no ground truth annotations are provided for the test set, we restricted our analysis to the annotated training data. The training images include annotations for the following object classes: car, van, truck, pedestrian, person (sitting), cyclist, tram and misc. In addition, some regions are labeled as ”don’t care” to mark areas with unlabeled objects, reducing the likelihood of false positives during evaluation.
For our case, we randomly split the annotated training set into training (80%) and validation (20%) subsets. To ensure representative coverage, we repeated the sampling process until approximately 80% of all ground truth pedestrian annotations were assigned to the training set and 20% to the validation set. The resulting dataset consists of 5,984 training images with 3,591 pedestrian objects and 1,497 validation images with 896 pedestrians. The object detectors were trained on all annotated classes, while evaluation was limited to the class pedestrian on the validation set. To infer predictions for all regions of an image, we did not consider don’t care regions during training and inference to not restrict the object detection task.
@@ IV-B Validated Ground Truth Generation
To establish a reliable evaluation benchmark, we constructed a high-quality validated ground truth (VGT) that achieves a very high recall by capturing every visible pedestrian. Each bounding box is associated with a probability score that reflects the likelihood of containing a pedestrian, rather than relying on a single hard label that may be erroneous. We use a two-stage microtask pipeline. In the first stage, we localize pedestrian instances. In the second stage, we validate the semantic content of each bounding box, generating soft labels that reflect annotator’s uncertainty.
| # Bounding Boxes | # Aggregations | Costs [s] | |
| Direct box | 2833 | 1.0 | 44,11 |
| Keypoint-to-box | 2099 | 3.0 | 92,671 |
| Combined box | 3078 | 2.4 | 103,79 |
@@
Firstly, our aim was to identify all possible human instances in the data. To achieve this, we used two complementary task configurations, each with three annotators.
Direct box annotation (Microtask 1): Multiple annotators draw tight bounding boxes around all the real humans that have not yet been marked in the image.
Keypoint-to-box annotation (Microtasks 2 & 3): Multiple annotators first place keypoints on unmarked humans, then draw a bounding box around each group of keypoints. The drawn bounding boxes are then aggregated.
Running both strategies in separate task configurations allows us to compare annotation effort and resulting recall. For constructing the VGT, we combined both methods to ensure maximal coverage of pedestrian instances. We manually removed duplicate boxes resulting from the different strategies. Further details and visual representations of the tasks are provided in Table I and the appendix.
Stage 2: Semantic validation.
In the second stage, we determine whether each bounding box contains a pedestrian. This process also serves as the foundation for error correction, as described in subsection III-C. For this purpose, we used three microtasks, each with 11 annotators (see Appendix for visual representations):
@@
While microtask 4 can theoretically provide sufficient information, the combination of Microtask 5 and 6 leads to increased annotator focus and reduced ambiguity. However, this approach requires twice the number of responses, as both tasks must be completed, resulting in 22 annotations per box.
@@
To further improve soft label quality, we incorporated two additional refinement steps. First, for boxes with high disagreement in microtasks 5 and 6, we added 11 additional annotations per task. This step, referred to as ambiguity refinement (AR). Second, we group duplicates of the same object from stage 1. These duplicate boxes were treated as describing the same object and aggregated accordingly. This grouping was applied exclusively to the VGT. An overview is given in Table II.
The final soft label for determining whether a bounding box contains a pedestrian is computed by multiplying the soft label from microtask 5 (real human) with the soft label from microtask 6 (standing or walking).
IV-C Object Detectors
The first step in our framework is to train the YOLOX and Cascade R-CNN object detectors on the KITTI dataset. For this, we used the OpenMMLab Detection Toolbox [63]. YOLOX was trained for 80 epochs on eight classes (car, van, pedestrian, cyclist, truck, misc, tram and person sitting). We chose a learning rate of 0.01, processing the images with a resolution of (640, 640) and a batch size of eight. On the test dataset, we performed evaluations and achieved an AP@50 of 77.8% for the class pedestrian with a score threshold of 0.3. To train Cascade R-CNN, we performed 70,000 iterations with a learning rate of 0.1. Images were batched with a batch size of six and have a resolution of (1000, 600). The trained model achieved an AP@50 of 63.1% for pedestrian with a score threshold of 0.3. After training, we made predictions on the test dataset filtering for the class pedestrian and a score threshold of 0.01. After applying non-maximum-suppression (NMS) with an threshold of 0.5, the remaining predictions were collected and form the basis of the label error detection methods.
IV-D Label Error Detection
We use the three label error detection methods introduced above, which are based on the predictions of the object detectors YOLOX and Cascade R-CNN and generate proposals for potential errors in the pedestrian annotations. We describe the implementation details of each method for generating bounding box proposals and the number of proposals to be reviewed. We also include the baseline of proposing detected bounding boxes ranked by objectness score.

### v2

@@ Abstract
Object detection has advanced rapidly in recent years, driven by increasingly large and diverse datasets. However, label errors often compromise the quality of these datasets and affect the outcomes of training and benchmark evaluations. Although label error detection methods for object detection datasets now exist, they are typically validated only on synthetic benchmarks or via limited manual inspection. How to correct such errors systematically and at scale remains an open problem. We introduce a semi-automated framework for label error correction called RECD (Rechecked). Building on existing label error detection methods, their error proposals are reviewed with lightweight, crowd-sourced microtasks. We apply RECD to the class pedestrian in the KITTI dataset, for which we crowdsourced high-quality corrected annotations. We detect 18% of missing and inaccurate labels in the original ground truth. We show that current label error detection methods, when combined with our correction framework, can recover hundreds of errors with little human effort compared to annotation from scratch. However, even the best methods still miss up to 66% of the label errors, which motivates further research, now enabled by our released benchmark.
@@ I Introduction
Deep neural networks (DNNs) have become state-of-the-art for extracting information from large-scale visual data [1, 2, 3, 4, 5, 6]. Training them for specialized and safety-critical domains, such as autonomous driving [1, 2, 5] or medical imaging [3, 6, 7, 8], requires vast amounts of high-quality annotated data [9, 10, 11, 12]. Crowdsourcing offers a cost- and time-effective solution for building large-scale vision datasets [13] and enables scalable annotation [14, 15, 16], particularly when tasks are broken down into simple, quick interactions known as microtasks [14, 13]. The quality of labels is critical because label errors such as missing labels, incorrect classes, or poor localization degrade model performance, generalization capability [17, 18, 19, 20, 21, 22, 23, 24] and benchmark validity [25, 17]. Although recent research has proposed automated methods for detecting such errors in image classification [25, 26, 27], semantic segmentation [28], and object detection [29], few works have focused on correcting them [26, 30, 31, 32] and none use error detection methods to reduce the re-annotation costs.
We propose RECD (see figure 1), a semi-automated framework that transitions from label error detection to correction for object detection datasets. RECD combines automated label error proposals from existing methods [33, 34, 35] with lightweight validation via microtasks. Aggregated responses capture both label correctness and ambiguity, enabling scalable and cost-effective correction. We apply RECD to the class pedestrian in the KITTI dataset [36]. The evaluation is intentionally limited to one class in order to establish proof of concept and control variables during the early stages of benchmarking. Our annotations reveal at least 18% of previously unlabeled or incorrectly located pedestrian labels. These corrected annotations will be made publicly available as a new benchmark for evaluating methods for detecting and correcting label errors. Our analysis shows that current detection methods can be rechecked faster than humans can draw the bounding boxes for labeling. However, up to 66% of errors remain undetected, highlighting the need for improved label error detection strategies, which is now supported by our benchmark.
@@
We propose RECD, a scalable and cost-efficient framework that utilizes state-of-the-art label error detection methods and extends them to label error correction in object detection datasets, enabling the correction of hundreds of errors faster than humans can annotate the bounding boxes.
We introduce ambiguity-aware labeling through soft labels yielding high quality object labels. We demonstrate that lower quality data might introduce more errors than it is fixing. We release corrected pedestrian annotations for KITTI, enabling the evaluation of label error detection methods on real-world label noise.
With our framework, we detected at least 18% of missing and inaccurate pedestrian annotations. However, we also find that up to 66% of the errors were not identified by current label error correction methods, which highlights the importance of this research, now enabled by our benchmark.
Our source code is publicly available at
https://github.com/JonathanKlees/rechecked.
@@ II Related Work
Identifying Label Errors in Object Detection Datasets. In recent years, there has been a research focus on label error detection in object detection. Cleanlab’s tool ObjectLab [35] assesses the quality of labels in object detection datasets by comparing predicted bounding boxes with GT, and provides a score describing the quality of the labels on an image. Another method, based on the instance-wise object detection loss [34], uses a two-stage object detector to detect real label errors by monitoring the regression and classification losses and using the discrepancy between predictions and ground truth labels. MetaDetect [33] is a method that performs meta classification and meta regression for predictions. Potential label errors are those predicted boxes that are evaluated as false positive ( with GT), but the meta classifier estimates a high probability for the prediction being correct. The method LidarMetaDetect [39] is an extension of MetaDetect that detects label errors in 3D object detection datasets. [40] developed a label refinement network that improves noisy localization labels. While we build on this recent research, we go one step further and aim at correcting these labels instead of just detecting them.
@@
Label Error Correction. In addition to detecting label errors, correcting them has also become an important step in improving label quality. [41] developed an active label cleaning framework to prioritize samples for re-annotation by comparing estimated label correctness and label difficulty. CROWDLAB [42] is another tool of Cleanlab that classifies crowdsourced labels for refinement to identify, which samples require re-annotation. In the work of [31], the datasets COCO [43] and Google Open Images [44] were re-annotated to compare model performance on original and re-annotated data. [32] combine foundation models with label corrections provided by human annotators to create PASCAL+, an updated version of the Pascal VOC dataset for semantic segmentation and[45] applied their re-labeling strategy, ReLabel, to re-annotate the ImageNet training set and published a corrected version. Our work focuses on the process of label error correction based on error detection, with the aim of reducing annotation costs.
@@
Ambiguity and Soft Labels in Computer Vision Datasets. In common computer vision benchmark datasets, images or objects are usually associated with a single hard class label. However, recent research shows that this concept is not representative of the ambiguity in real data [46, 47, 48, 49, 50, 51, 52, 53, 54, 23] and even harms model performance [20, 21, 22, 55, 23, 24]. In contrast to this, soft or probabilistic labels [56] indicate a probability distribution across several classes, reflecting that the correct label for an instance is not always straightforward and multiple annotators may disagree on the label. Soft labels in object detection datasets are rare but there exist methods that simulate soft labels based on single-class annotations through deep learning [57, 58, 59]. We extend this research by showing that the label quality is essential for correcting labels and poor quality data might even introduce more errors than it corrects. Furthermore, we release these soft labels as a benchmark for future label error detection and correction research.
@@ III Method RECD
The most challenging task in label error detection is detecting missing labels, i.e., objects present in the image that are missing in the GT annotations. Unlike inaccurate bounding box localizations and incorrect class labels, which only require verifying existing annotations and consequently, are easier to identify, detecting missing labels requires a complete review of each image. We propose a semi-automatic framework that combines label error detection methods with a microtask-based review process to efficiently correct label errors. As depicted in figure 2, our framework relies on the predictions of object detectors, which are then evaluated through label error detection methods resulting in a much smaller number of label error proposals to be manually verified and corrected.
III-A Label Error Detection Methods
RECD is model-agnostic. Any object detector can be used to infer initial predictions. In our experiments, we use YOLOX [60], a one-stage detector, and Cascade R-CNN [61], a two-stage detector. We combine them with existing label error detection methods including MetaDetect [33], loss-based instance-wise scoring [34] and ObjectLab [35]. The considered methods compare the predictions of an object detector with the GT annotations to identify overlooked objects or misfitting bounding boxes. The predictions are then scored by the label error detection method, reflecting the likelihood of a label error being present. This way, we obtain label error proposals, which are then reviewed manually in the second stage.
III-B Error Correction
From the previous stage, we obtain a collection of label error proposals that need to be manually evaluated. We created short and easy-to-answer questions consisting of a single visual task, e.g., whether the highlighted object is a human or not, as illustrated in figure 3. These kinds of micro-questions are also referred to as microtasks [62]. With microtasks, annotators can quickly answer one simple question at a time. We repeat the microtasks with several people, which results in an estimate of the underlying distribution of the outcomes associated to this question. More complex questions, such as ’What is a pedestrian?’, are broken down into multiple microtasks, such as ’Is it a human?’ and ’What is the person doing?’. See IV-B for the microtasks used to construct the validated annotations of the KITTI dataset. We can use the calculated soft label probability to determine whether a proposed label error is indeed erroneous or not.
@@ IV-A Dataset Construction
We apply RECD to the KITTI [36] 2D object detection dataset. We focus on the class pedestrian as a challenging and safety-relevant use case. The 7,481 training images include annotations for the classes: car, van, truck, pedestrian, person (sitting), cyclist, tram and misc. In addition, some regions are labeled as ’don’t care’ to mark areas with unlabeled objects. Since no GT annotations are provided for the test set, we randomly split the annotated training set into training (80%) and validation (20%) subsets. To ensure representative coverage, we repeated the sampling process until approximately 80% of all GT pedestrian annotations were assigned to the training set and 20% to the validation set. The resulting dataset consists of 5,984 training images with 3,591 pedestrian objects and 1,497 validation images with 896 pedestrian annotations.
| # Bounding Boxes | # Aggregations | Costs [s] | |
| Direct box | 2833 | 1.0 | 44.11 |
| Keypoint-to-box | 2099 | 3.0 | 92.67 |
| Combined box | 3078 | 2.4 | 103.79 |
@@ IV-B Validated Ground Truth Generation
To establish a reliable evaluation benchmark, we constructed a high-quality validated ground truth (VGT) that achieves a very high recall by capturing every visible pedestrian. Each bounding box is associated with a probability score that reflects the likelihood of containing a pedestrian, rather than relying on a single hard label that may be erroneous.
We use a two-stage microtask pipeline. In the first stage, we localize pedestrian instances. In the second stage, we validate the semantic content of each bounding box, generating soft labels that reflect annotator’s uncertainty.
@@
Our first aim is to identify all possible human instances in the data. To achieve this, we used two complementary task configurations, each with three different annotators.
Direct box annotation (Microtask 1): All annotators draw tight bounding boxes around humans that have not yet been marked in the image.
Keypoint-to-box annotation (Microtasks 2 & 3): Each annotator first places a single keypoint on unmarked humans. Then, around each resulting group of keypoints, each annotator draws a bounding box. The drawn bounding boxes are then aggregated.
Running both strategies in separate task configurations allows us to compare annotation effort and resulting recall. For constructing the VGT, we combined both methods to ensure maximal coverage of pedestrian instances. We manually removed duplicate boxes resulting from the different strategies. Further details and visual representations of the tasks are provided in table I and figure 4.
@@
Stage 2: Semantic validation.
In the second stage, we determine whether each bounding box contains a pedestrian. This process also serves as the foundation for label error correction, as described in III-B. For this purpose, we used the following three microtasks, each of which was answered by 11 different annotators:
@@
While microtask 4 can theoretically provide sufficient information, the combination of microtask 5 and 6 leads to increased annotator focus and reduced ambiguity. However, this approach requires twice the number of responses, as both tasks must be completed, resulting in 22 annotations per bounding box. To further improve soft label quality, we incorporated two additional refinement steps. First, for boxes with high disagreement in microtasks 5 and 6, we added 11 additional annotations per task. We refer to this step as ambiguity refinement (AR). Second, we group duplicates of the same object from stage 1. These duplicate boxes were treated as describing the same object and aggregated accordingly. This grouping was applied exclusively to the VGT. An overview is given in table II. The final soft label for determining whether a bounding box contains a pedestrian is computed by multiplying the soft label from microtask 5 (real human) with the soft label from microtask 6 (standing or walking).
IV-C Label Error Detection
We use three label error detection methods, which are based on the predictions of YOLOX and Cascade R-CNN. These methods generate proposals for potential label errors.